# The Codecfake Dataset and Countermeasures for the Universally Detection of Deepfake Audio

### v2

@@ Abstract
With the proliferation of Audio Language Model (ALM) based deepfake audio, there is an urgent need for generalized detection methods. ALM-based deepfake audio currently exhibits widespread, high deception, and type versatility, posing a significant challenge to current audio deepfake detection (ADD) models trained solely on vocoded data. To effectively detect ALM-based deepfake audio, we focus on the mechanism of the ALM-based audio generation method, the conversion from neural codec to waveform. We initially construct the Codecfake dataset, an open-source large-scale dataset, including 2 languages, over 1M audio samples, and various test conditions, focus on ALM-based audio detection. As countermeasure, to achieve universal detection of deepfake audio and tackle domain ascent bias issue of original SAM, we propose the CSAM strategy to learn a domain balanced and generalized minima. In our experiments, we first demonstrate that ADD model training with the Codecfake dataset can effectively detects ALM-based audio. Furthermore, our proposed generalization countermeasure yields the lowest average Equal Error Rate (EER) of 0.616% across all test conditions compared to baseline models. The dataset [^1] https://zenodo.org/records/11171708and associated code [^2] https://github.com/xieyuankun/Codecfakeare available online.
@@ I Introduction
In this paper, we propose Codecfake, an initial large-scale dataset for detecting ALM-based deepfake audio, encompassing over 1M of audio samples across two languages. Specifically, we select 7 representative open-source neural codec models to construct fake audio. These codec types encompass the current mainstream ALM-based audio generation models. To effectively assess the capability of ADD models trained with Codecfake, we set up diverse test conditions, encompassing unseen codec-based audio, codec-based audio synthesized with varying parameters, and in-the-wild ALM-based audio. For countermeasure, we propose a generalized training strategy called co-training sharpnes aware minimization (CSAM) to effective detect both the vocoder-based audio and codec-based audio. Experiments demonstrate that vocoder-trained ADD models are insufficient for effectively detecting codec-based audio. Leveraging the proposed Codecfake dataset and CSAM strategy for training, we achieve the lowest 0.616% average EER on diverse test condition.
@@ II-C Audio Deepfake Detection Dataset
With the fast propagation and high-quality of deepfake audio, the filed of ADD has attracted increasing attention from endeavors. The first step in expanding a field of research is to establish corresponding datasets. In the current task of ADD, numerous dataset initiatives are propelling the development of the field.
@@
Surrounding the ASVspoof Challenge series, especially in ASVspoof2019LA (19LA)[2], numerous outstanding works on ADD have emerged. The training set of 19LA includes six vocoder-based synthesized TTS and VC utterances, labeled as A01-A06. The test set of 19LA comprises unseen spoofing methods A07-A19. This training and testing protocol aims to select generalizable anti-spoofing models. In subsequent challenges like ASVspoof2021LA (21LA) and ASVspoof2021DF (21DF) track [3], participants are still limited to using the training set from 19LA. However, the test sets introduce more challenging conditions. Specifically, 21LA track includes 19LA evaluation set condition from different real telephone systems, while the 21DF track involves more than hundreds of vocoder-based audio samples generated from various source domains and under various compression conditions.
@@
In addition to the two challenge series datasets mentioned, there are also other ADD datasets designed to address various issues. Wavefake [55] presents a vocoder-based audio deepfake dataset collected from ten sample sets across six different network architectures, spanning two languages. In the wild (ITW) dataset [7] collects both real and fake audio samples from publicly available sources like social networks and video streaming platforms. It potentially contains background noises and is designed to assess the generalizability of ADD models. CFAD [56] is the first Mandarin audio spoofed dataset consisting of twelve different vocoder-based generation algorithms including both TTS and VC types. Although current ADD datasets offer a rich variety of fake types, most of them consist of vocoder-based audio. With the advancements in ALM generation method, there is an urgent need for datasets generated based on codecs to facilitate research into detecting ALM-based audio.
@@ III-A Overview
Motivation. We employ speech codec models to construct a dataset of forged audio, aiming to enhance the detection performance of fake audio detection models against ALM-based audio. This choice is motivated by the fact that, as illustrated in Figure 1, existing ALMs universally adopt neural audio codecs to construct discrete encodings. These models utilize their respective decoders to transform tokens generated by ALMs into raw audio. Consequently, the audio produced by various ALMs contains artifacts generated by the neural audio codecs decoder. By discerning these artifacts, which arise from the reconstruction of audio by the neural audio codecs, one can effectively determine whether the audio has been forged by ALMs. On the other hand, from a pragmatic standpoint, the majority of current ALMs are not open-source, and the cost of training is prohibitive for most individuals. In contrast, neural audio codec models often have reliable implementations and are widely utilized by ALMs.
Difference. Traditional ADD datasets are primarily constructed using vocoders, while ALM primarily utilizes codec-based methods to decode discrete codes generated by language models into audio, rather than using vocoders. Consequently, models trained on vocoder-based datasets have limited generalization capabilities for detecting ALM-generated fake audio. To address this limitation, this dataset employs various neural audio codec models to construct the ADD dataset. Additionally, during training, audio codec models often include data beyond speech, such as general audio, vocals, music, etc., which differs from traditional vocoders that are trained solely on speech. Moreover, audio codec models typically do not use handcraft acoustic features as intermediate features during training, which is a notable distinction from traditional vocoders that utilize handcraft acoustic features. Training ADD models using a combination of codec-based audio and audio synthesized by traditional vocoders can significantly enhance model generalization. To validate the detection performance of models trained on codec-based audio data for ALM-based audio, we included language-based audio data in the Codecfake Dataset.
@@ III-B Architectures for generating codec-based fake audio
TABLE I: Neural codec models detail.
@@
TABLE II: Different parameters setting for test condition C3 and C4.
@@ III-C The Generation Process of codec-based fake audio
During the inference phase, we utilize the trained seven neural codec models to re-encode and decode data from the VCTK [63] (This CSTR VCTK Corpus includes speech data uttered by 110 English speakers with various accents.) and AISHELL3 [64] (The corpus contains roughly 85 hours of emotion-neutral recordings spoken by 218 native Chinese mandarin speakers) datasets separately. More neural codec models details is ilustrated in Table I.
@@ III-E Testing Condition
To verify whether codec-trained ADD model can detect audio generated by ALM, we created ALM conditions A1-A3. The generation of these three tasks adopts their open-source default configurations. Among them, A1 and A2 represent text-to-speech (TTS) task, while A3 represents a text-to-audio (TTA) task. A1 is generated by VALL-E [34], A2 by VALL-E X [36], and A3 by AudioGen [33]. For A1 and A2, we selct the text sourced from the VCTK evaluation set corresponding to 4451 text items. However, the speaker p315 does not have corresponding text, so the generated audio for A1 and A2 consists of only 4437 audio samples. For the real source domain of A3, we utilize the training subset of the Audiocaps [65], which was employed during the training phase of AudioGen, generating 49,838 fake audios based on its corresponding texts. Given that some of the original Audiocaps audio links are no longer active, we use the currently available 49,274 audio samples as the real source domain for A3.
@@ IV-A Baseline model
For ADD model, we select three baseline models for evaluation as shown in Figure 4. For the front-end features, we chose the commonly used handcrafted feature in the field of audio deepfake generation, the Mel-spectrogram, to facilitate the analysis of different codec-based audios. On the other hand, we utilized the frozen pre-trained hidden representation of wav2vec2-XLS-R (W2V2-XLS-R) [66] as features. This is because existing research has already convincingly demonstrated the superior performance of wav2vec2-xls-r in the ADD task [67]. The front-end W2V2-XLS-R is pre-trained with 436k hours of unannotated genuine speech data in 128 languages using a contrastive objective. As a result of pre-training, the hidden states from the transformer can represent the genuine speech contextualized information. It is worth mentioning that we utilize the fifth hidden state from the transformer as the final feature, in reference to previous research on frozen hidden states [68], which demonstrated the best performance.
@@
Figure 4: Three baseline ADD model for evaluation. (a) represents Mel-LCNN, (b) represents W2V2-LCNN, and (c) represents W2V2-AASIST.
@@ V-A Experiments setting
At first, to validate whether training the ADD model using conventional vocoder-based audio can effectively detect codec-based audio, we construct the vocoder-trained ADD model using the training set of ASVspoof2019LA (19LA). 19LA was created utilizing 107 speakers (46 male, 61 female) utterances and contains a total number of 12,456 real and 108,978 fake speeches by using 11 TTS and 8 VC (A01-A19) spoofing algorithms. The training and development sets consist of data from spoofing systems denoted by A01 to A06, while the evaluation set includes attacks from spoofing systems denoted by A7 to A19. The details of 19LA can be found in Table V.
TABLE III: EER (%) results for ADD methods trained by 19LA training set. CAVG represents the average EER across C1-C7.
Model 19LA IWA C1 C2 C3 C4 C5 C6 C7 CAVG AVG
TABLE IV: EER (%) results for ADD methods trained by Codecfake training set. CAVG represents the average EER across C1-C7.
Model 19LA IWA C1 C2 C3 C4 C5 C6 C7 CAVG AVG
TABLE V: Summary of the 19LA dataset.
@@ V-B Implementation details
For pre-processing of the ADD baseline models, all audio samples were first down-sampled to 16,000 Hz and trimmed or padded to a duration of 4 seconds. For the mel-spectrogram, we extracted a 80-dimensional mel-spectrogram. For W2V2, we employed the Wav2Vec-XLS-R [^3] https://huggingface.co/facebook/wav2vec2-xls-r-300m model with frozen parameters, extracting the 1024-dimensional hidden states as the feature representation.
@@ VI-A Vocoder-trained ADD model
First, we explore whether the vocoder-trained ADD model can effectively detect codec-based audio, as outlined in Table III. Concretely, we conducted training on three baseline models: Mel-LCNN, W2V2-LCNN and W2V2-AASIST, utilizing the ASVspoof2019LA (19LA) [74] training set. From the results, we can first observe that they achieve promising result in 19LA test sets and W2V2-AASIST achieve the lowest EER with 0.122%. As the 19LA training set incorporates six spoofing methods, and the test set encompasses a total of unseen 13 methods, achieving favorable results on the 19LA test set suggests that the detection system exhibits generalizability. However, when the vocoder-trained ADD models evaluate on the Codecfake test sets, the results are quite unfavorable. Mel-LCNN, W2V2-LCNN, and W2V2-AASIST achieve average EER of 39.941%, 32.764%, and 34.991%, respectively, which represents a decrease in performance compared to the 19LA test set. This indicates that models trained with vocoders cannot generalize to effectively detect codec-based audio.
@@
TABLE VI: EER (%) results for ADD methods trained by Codecfake training set and test on different parameter setting on C3 and C4.
@@
TABLE VII: EER (%) results for ADD methods test on ALM-based audio.
@@
VI-B Codec-trained ADD model
Due to the poor performance of vocoder-trained ADD models on the Codecfake test set, we proceed to train the baseline models using the Codecfake training set. The results are presented in the Table IV. Mel-LCNN, W2V2-LCNN, and W2V2-AASIST achieve average EER of 17.987%, 1.983%, and 1.628%, exhibited reductions in average EER of 21.954%, 30.781%, and 33.363%, respectively, compared to the vocoder-trained model. All ADD models trained with the Codecfake dataset exhibit promising results under Codecfake test condition C1-C7. Among these, W2V2-AASIST achieved the best average in EER of 0.177%. In the unseen codec test condition C7, W2V2-AASIST achieves an EER of 0.884%, indicating that training with the Codecfake dataset can generalize to detecting unknown codec-based audio. Furthermore, we found that using only the codec-trained ADD model can detect vocoder-based audio. Specifically, W2V2-AASIST achieves a 3.806% EER on the 19LA test set, and W2V2-LCNN achieves a 4.975% EER on the ITW dataset. For the ITW results, as far as we know, they already surpass all the ADD models trained on vocoder-based audio.

### v3

@@ Abstract
With the proliferation of Audio Language Model (ALM) based deepfake audio, there is an urgent need for generalized detection methods. ALM-based deepfake audio currently exhibits widespread, high deception, and type versatility, posing a significant challenge to current audio deepfake detection (ADD) models trained solely on vocoded data. To effectively detect ALM-based deepfake audio, we focus on the mechanism of the ALM-based audio generation method, the conversion from neural codec to waveform. We initially constructed the Codecfake dataset, an open-source, large-scale collection comprising over 1 million audio samples in both English and Chinese, focus on ALM-based audio detection. As countermeasure, to achieve universal detection of deepfake audio and tackle domain ascent bias issue of original sharpness aware minimization (SAM), we propose the CSAM strategy to learn a domain balanced and generalized minima. In our experiments, we first demonstrate that ADD model training with the Codecfake dataset can effectively detects ALM-based audio. Furthermore, our proposed generalization countermeasure yields the lowest average equal error rate (EER) of 0.616% across all test conditions compared to baseline models. The dataset [^1] https://zenodo.org/records/13838106and associated code [^2] https://github.com/xieyuankun/Codecfakeare available online.
@@ I Introduction
In this paper, we introduce Codecfake, a initial large-scale dataset for detecting ALM-based deepfake audio, comprising over 1 million audio samples in two languages: English and Chinese. Specifically, we select 7 representative open-source neural codec models to construct fake audio. These codec types encompass the current mainstream ALM-based audio generation models. To effectively assess the capability of ADD models trained with Codecfake, we set up diverse test conditions, encompassing unseen codec-based audio, codec-based audio synthesized with varying parameters, and in-the-wild ALM-based audio. For countermeasure, we propose a generalized training strategy called co-training sharpness aware minimization (CSAM) to effective detect both the vocoder-based audio and codec-based audio. Experiments demonstrate that vocoder-trained ADD models are insufficient for effectively detecting codec-based audio. Leveraging the proposed Codecfake dataset and CSAM strategy for training, we achieve the lowest 0.616% average EER on diverse test condition.
@@ II-B Codec-based deepfake audio
TABLE I: Characteristics of representative datasets in the field of ADD.
Dataset Year Languages Methods Back-end Type Real Fake Total
ASVspoof2019LA [2] 2019 English 19 Vocoder Speech 10,256 90,192 100,448
ASVspoof2021LA [3] 2021 English 19 Vocoder Speech 14,816 133,360 148,176
ASVspoof2021DF [3] 2021 English 100+ Vocoder Speech 14,869 519,059 533,928
WaveFake [54] 2021 English/Japanese 7 Vocoder Speech 0 117,985 117,985
ITW [7] 2022 English - - Speech 19,963 11,816 31,779
TIMIT-TTS [55] 2022 English 12 Vocoder Speech 0 5,160 5,160
ADD2022T1.2 [4] 2022 Chinese - Vocoder Speech 36,953 123,932 160,885
CFAD [56] 2023 Chinese 12 Vocoder Speech 38,600 77,200 115,800
ADD2023T1.2 [57] 2023 Chinese - Vocoder Speech 172,819 113,042 285,861
MLAAD [58] 2024 Multilanguages 54 Vocoder Speech 0 76,000 76,000
ASVspoof5 [59] 2024 English 32 Vocoder Speech 188,819 815,262 1,004,081
Codecfake 2024 English/Chinese 7 Neural Codec Speech/Sound 132,277 925,939 1,058,216
@@ II-C Audio Deepfake Detection Dataset
With the fast propagation and high-quality of deepfake audio, the filed of ADD has attracted increasing attention from endeavors. The first step in expanding a field of research is to establish corresponding datasets. In the current task of ADD, numerous dataset initiatives are propelling the development of the field. In the Table I,
@@
Surrounding the ASVspoof Challenge series, especially in ASVspoof2019LA (19LA)[2], numerous outstanding works on ADD have emerged. The training set of 19LA includes six vocoder-based synthesized TTS and VC utterances, labeled as A01-A06. The test set of 19LA comprises unseen spoofing methods A07-A19. This training and testing protocol aims to select generalizable anti-spoofing models. In subsequent challenges like ASVspoof2021LA (21LA) and ASVspoof2021DF (21DF) track [3], participants are still limited to using the training set from 19LA. However, the test sets introduce more challenging conditions. Specifically, 21LA track includes 19LA evaluation set condition from different real telephone systems, while the 21DF track involves more than hundreds of vocoder-based audio samples generated from various source domains and under various compression conditions. Currently, ASVspoof5 [59] has been successfully held, focusing on crowdsourced datasets and including adversarial attacks and various compression conditions.
@@
In addition to the two challenge series datasets mentioned, there are also other ADD datasets designed to address various issues. WaveFake [54] presents a vocoder-based audio deepfake dataset collected from ten sample sets across six different network architectures, spanning two languages. In the wild (ITW) dataset [7] collects both real and fake audio samples from publicly available sources like social networks and video streaming platforms. It potentially contains background noises and is designed to assess the generalizability of ADD models. CFAD [56] is the first Mandarin audio spoofed dataset consisting of twelve different vocoder-based generation algorithms including both TTS and VC types. MLAAD [58] is a recent work that covers 23 languages and up to 54 TTS and VC spoofing methods.
Although current ADD datasets offer a rich variety of spoof types, most of their generation back-ends are based on vocoders. With the advancements in ALM generation method, there is an urgent need for datasets generated based on codecs to facilitate research into detecting ALM-based audio.
@@ III-A Overview
Motivation. We employ speech codec models to construct a dataset of forged audio, aiming to enhance the detection performance of fake audio detection models against ALM-based audio. This choice is motivated by:
  • Improving generalizability. Traditional ADD datasets are primarily constructed using vocoders, while ALM primarily utilizes codec-based methods to decode discrete codes generated by language models into audio, rather than using vocoders. Consequently, models trained on vocoder-based datasets have limited generalization capabilities for detecting ALM-generated fake audio.
  • Strengthening codec representation learning. ALM-based audio decoders apply neural codec algorithms to convert discrete codes into waveforms, which can often give rise to artifacts due to the upsampling convolution module[60]. Training with Codecfake enables ADD models to learn the patterns of codec artifacts, thus enhancing their ability to effectively detect ALM-based deepfake audio.
  • Practical needs. From a practical standpoint, most current ALMs are not open-source, and the cost of training them is prohibitive for most individuals. In contrast, neural audio codec models have reliable implementations and are widely utilized in ALM systems.
Strength. Compared to existing ADD datasets, our proposed Codecfake dataset presents several key advantages:
  • As shown in Table I, traditional ADD datasets are primarily constructed using vocoders, whereas our proposed Codecfake dataset leverages neural codec-based audio, enabling more effective detection of ALM-based audio.
  • The Codecfake dataset provides a wide range of test conditions simulating real-world scenarios, encompassing various audio types (speech and sound) and ALM-based audio (A1, A2).
  • Our proposed Codecfake dataset comprises over 1 million samples across two languages and covers codecs used in current ALM implementations.
@@ III-B Architectures for generating codec-based fake audio
TABLE II: Neural codec models detail.
@@
TABLE III: Different parameters setting for test condition C3 and C4.
@@ III-C The Generation Process of codec-based fake audio
During the inference phase, we utilize the trained seven neural codec models to re-encode and decode data from the VCTK [67] (This CSTR VCTK Corpus includes speech data uttered by 110 English speakers with various accents.) and AISHELL3 [68] (The corpus contains roughly 85 hours of emotion-neutral recordings spoken by 218 native Chinese mandarin speakers) datasets separately. More neural codec models details is ilustrated in Table II.
@@ III-E Testing Condition
To verify whether codec-trained ADD model can detect audio generated by ALM, we created ALM conditions A1-A3. The generation of these three tasks adopts their open-source default configurations. Among them, A1 and A2 represent text-to-speech (TTS) task, while A3 represents a text-to-audio (TTA) task. A1 is generated by VALL-E [34], A2 by VALL-E X [36], and A3 by AudioGen [33]. For A1 and A2, we select the text sourced from the VCTK evaluation set corresponding to 4451 text items. However, the speaker p315 does not have corresponding text, so the generated audio for A1 and A2 consists of only 4437 audio samples. For the real source domain of A3, we utilize the training subset of the Audiocaps [69], which was employed during the training phase of AudioGen, generating 49,838 fake audios based on its corresponding texts. Given that some of the original Audiocaps audio links are no longer active, we use the currently available 49,274 audio samples as the real source domain for A3.
@@ IV-A Baseline model
For ADD model, we select four baseline models for evaluation as shown in Figure 4. For the front-end features, we chose the commonly used handcrafted feature in the field of audio deepfake generation, the Mel-spectrogram, to facilitate the analysis of different codec-based audios. On the other hand, we utilized the frozen pre-trained hidden representation of Wav2vec2-XLS-R (W2V2) [70] and WavLM [71] as input features. For W2V2, existing research has convincingly demonstrated its superior performance in the ADD task [72, 73, 74]. Besides W2V2, we also adopt WavLM as the front-end feature due to its exceptional performance in other downstream tasks [75]. It is worth mentioning that we utilize the fifth hidden state from the frozen pre-trained feature as the final feature, in reference to previous research on frozen hidden states [76], which demonstrated the best performance.
@@
Figure 4: Four baseline ADD models for evaluation. (a), (b), (c), (d) represents Mel-LCNN, W2V2-LCNN, WavLM-AASIST, W2V2-AASIST, respectively.
@@ V-A Experiments setting
At first, to validate whether training the ADD model using conventional vocoder-based audio can effectively detect codec-based audio, we construct the vocoder-trained ADD model using the training set of ASVspoof2019LA (19LA). 19LA was created utilizing 107 speakers (46 male, 61 female) utterances and contains a total number of 12,456 real and 108,978 fake speeches by using 11 TTS and 8 VC (A01-A19) spoofing algorithms. The training and development sets consist of data from spoofing systems denoted by A01 to A06, while the evaluation set includes attacks from spoofing systems denoted by A7 to A19. The details of 19LA can be found in Table IV.
TABLE IV: Summary of the 19LA dataset.
@@
TABLE V: EER (%) results for ADD methods trained by 19LA training set. CAVG represents the average EER across C1-C7.
Model 19LA ITW C1 C2 C3 C4 C5 C6 C7 CAVG AVG
WavLM-AASIST 0.462 17.917 41.752 53.122 50.469 39.205 44.224 48.397 47.838 46.430 38.154
TABLE VI: EER (%) results for ADD methods trained by Codecfake training set. CAVG represents the average EER across C1-C7.
Model 19LA ITW C1 C2 C3 C4 C5 C6 C7 CAVG AVG
WavLM-AASIST 14.455 22.528 0.045 0.121 0.839 0.098 0.076 3.818 15.270 2.895 6.361
@@ V-B Implementation details
For pre-processing of the ADD baseline models, all audio samples were first down-sampled to 16,000 Hz and trimmed or padded to a duration of 4 seconds. For the mel-spectrogram, we extracted a 80-dimensional mel-spectrogram. For self-supervised pre-trained feature, we employed the Wav2Vec-XLS-R [^3] https://huggingface.co/facebook/wav2vec2-xls-r-300m and WavLM-large [^4] https://huggingface.co/microsoft/wavlm-large with frozen parameters, extracting the 1024-dimensional hidden states as the feature representation.
@@ VI-A Vocoder-trained ADD model
First, we explore whether the vocoder-trained ADD model can effectively detect codec-based audio, as outlined in Table V. Concretely, we conducted training on four baseline models: Mel-LCNN, W2V2-LCNN, WavLM-AASIST and W2V2-AASIST, utilizing the ASVspoof2019LA (19LA) [82] training set. From the results, we can first observe that they achieve promising result in 19LA test sets and W2V2-AASIST achieve the lowest EER with 0.122%. As the 19LA training set incorporates six spoofing methods, and the test set encompasses a total of unseen 13 methods, achieving favorable results on the 19LA test set suggests that the detection system exhibits generalizability. However, when the vocoder-trained ADD models evaluate on the Codecfake test sets, the results are quite unfavorable. Mel-LCNN, W2V2-LCNN, WavLM-AASIST and W2V2-AASIST achieve average EER of 39.941%, 32.764%, 38.154% and 34.991%, respectively, which represents a decrease in performance compared to the 19LA test set. This indicates that models trained with vocoders cannot generalize to effectively detect codec-based audio.
@@
VI-B Codec-trained ADD model
Due to the poor performance of vocoder-trained ADD models on the Codecfake test set, we proceed to train the baseline models using the Codecfake training set. The results are presented in the Table VI. Mel-LCNN, W2V2-LCNN, WavLM-AASIST and W2V2-AASIST achieve average EER of 17.987%, 1.983%, 6.361% and 1.628%, exhibited reductions in average EER of 21.954%, 30.781%, 31.793% and 33.363%, respectively, compared to the vocoder-trained model. All ADD models trained with the Codecfake dataset exhibit promising results under Codecfake test condition C1-C7. Among these, W2V2-AASIST achieved the best average in EER of 0.177%. In the unseen codec test condition C7, W2V2-AASIST achieves an EER of 0.884%, indicating that training with the Codecfake dataset can generalize to detecting unknown codec-based audio. Furthermore, we found that using only the codec-trained ADD model can detect vocoder-based audio. Specifically, W2V2-AASIST achieves a 3.806% EER on the 19LA test set, and W2V2-LCNN achieves a 4.975% EER on the ITW dataset. For the ITW results, as far as we know, they already surpass all the ADD models trained on vocoder-based audio.
TABLE VII: EER (%) results for ADD methods trained by Codecfake training set and test on different parameter setting on C3 and C4.
@@ VI-B Codec-trained ADD model
TABLE VIII: EER (%) results for the impact of language setting. All experiments use the top-performing ADD model, W2V2-AASIST. ’English’ refers to training or testing with the VCTK subset of Codecfake, while ’Chinese’ denotes the ALSHELL subset. ’ALL’ indicates the original multilingual testing conditions of Codecfake.
Train Test C1 C2 C3 C4 C5 C6 C7 AVG
English English 0.000 0.000 0.045 0.022 0.000 0.247 1.123 0.205
English Chinese 0.763 1.219 1.618 3.509 1.128 4.295 14.584 3.874
English ALL 0.302 0.605 1.225 1.875 0.514 4.445 10.425 2.770
Chinese English 0.135 0.764 3.595 0.247 0.247 8.627 14.626 4.034
Chinese Chinese 0.182 0.057 0.330 0.091 0.046 1.686 15.096 2.498
Chinese ALL 1.051 0.431 7.998 0.370 0.227 12.799 26.293 7.024
TABLE IX: EER (%) results for ADD methods test on ALM-based audio.